[Brahms-l] Shift report 20040131 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Sat Jan 31 2004 - 07:57:33 EST
------------------------------------------------------------
Date: Sat Jan 31 07:57:32 2004, Shift: 00:00-08:00
Supervisor: rk, with: 
Log book#15 pages 46 - 46

Shift Summary:
------------------------------------------------------------
Beam for about 3.5 hours, very poor BLUE - only few bunches
B6esides standard pedestal runs (9773, 9774) the following runs have been done:
9975 - tr.2 150K @ FS1723B6deg, tr.3 114K @ MRS2255B40deg
9976 - tr.2 182K @ FS1723B6deg, tr.3 182K @ MRS2255B40deg
9977 - tr.2  37K @ FS1723B6deg, tr.3  48K @ MRS2255B40deg
No detector problems were observed.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sat Jan 31 07:57:48 2004

This archive was generated by hypermail 2.1.8 : Sat Jan 31 2004 - 07:58:03 EST